OpenStudy (anonymous):

Find all real solutions. show work: \[2(x-4)^{7\over3} - (x-4)^{4\over3} - (x-4)^ {1\over3}=0\]

OpenStudy (anonymous):

anybody? i don't know where to start...

OpenStudy (anonymous):

4 is a solution automatically

OpenStudy (anonymous):

5 is the answer.... :/

OpenStudy (amistre64):

ack!! ... wolframalpha.com; :)

OpenStudy (anonymous):

i really need to know how to do it. lol. expected to be on my test...

OpenStudy (anonymous):

wolframalpha told me the answer is 5 ;)

OpenStudy (amistre64):

its the same as other polys; but they try to throw you with the rational exponents :)

OpenStudy (amistre64):

you can factor out a ()1/3 and solve the quadratic that is left over

OpenStudy (amistre64):

u = (x-4)^1/3 \[2u^7-u^4-u = 0\] \[u(2u^6-u^3-1) = 0\] \[u[2(u^3)^2-(u^3)-1] = 0\] this amounts to just solving for: sub again like maybe: a = u^3 2a^2 -a -1

OpenStudy (amistre64):

a = 2/2 +- sqrt(1-4(-1)(2))/2 = 1 +- sqrt(9)/2 = 1 +- 3/2; 2.5 and -.5 ................................................ u^3 = 2.5 and/or u^3 = -.5 u = cbrt(2.5) u = cbrt(-.5) ............................................... cbrt(x-4) = u cbrt(x-4) = cbrt(2.5) and/or cbrt(x-4) = cbrt(-.5) x-4 = 2.5 x-4 = -.5 x = 6.5 x = 3.5 I prolly missed it in there someplace but I blame the problem :)

OpenStudy (amistre64):

ack .. i used 2 instead of 1 in my quad formula

OpenStudy (anonymous):

A bit of a cheat solution, but you know that 4 is a solution since all the brackets are zero so sum to zero. Then since the coefficients sum to zero and the brackets are all the same. Make sure they equal 1 (since 1^n=1). So (x-4)=1 so x=5.

OpenStudy (amistre64):

the simple way is prolly best ;)